OpenStudy (anonymous):

\[f(x)=x ^{2}-8 \ln(x)\] where ln denotes the natural logarithm function. a)found intervals of increase and decrease b) calculates the absolute and relative extremes of the function f c) concavity and convexity

OpenStudy (anonymous):

f'(x) nikalo...

OpenStudy (anonymous):

is the problem

OpenStudy (anonymous):

f(x)= 2x-8/x then?

OpenStudy (anonymous):

set it <0 and >0

OpenStudy (anonymous):

hmmm k

OpenStudy (aravindg):

If f'(x)>0 for all x belongs to I , then f is increasing on I. If f'('x)<0 for all belongs to I, then f is decreasing on I.

OpenStudy (anonymous):

f(x)= 2x-8/x=0 how can find point

OpenStudy (anonymous):

f(x)= 2x-8/x=0 2x-8=x 2x-x-8=0 x-8=0 x=8 is correct??

OpenStudy (anonymous):

@Yahoo!

OpenStudy (nubeer):

no its not correct...

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

:)

OpenStudy (nubeer):

ok

OpenStudy (anonymous):

:)